# CDR-Agent: Intelligent Selection and Execution of Clinical Decision Rules Using Large Language Model Agents

### v1

@@ Abstract
Clinical decision-making is inherently complex and fast-paced, particularly in emergency departments (EDs) where critical, rapid and high-stakes decisions are made. Clinical Decision Rules (CDRs) are standardized evidence-based tools that combine signs, symptoms, and clinical variables into decision trees to make consistent and accurate diagnoses. CDR usage is often hindered by the clinician’s cognitive load, limiting their ability to quickly recall and apply the appropriate rules. We introduce CDR-Agent, a novel LLM-based system designed to enhance ED decision-making by autonomously identifying and applying the most appropriate CDRs based on unstructured clinical notes. To validate CDR-Agent, we curated two novel ED datasets: synthetic and CDR-Bench, although CDR-Agent is applicable to non ED clinics. CDR-Agent achieves a 56.3% (synthetic) and 8.7% (CDR-Bench) accuracy gain relative to the standalone LLM baseline in CDR selection. Moreover, CDR-Agent significantly reduces computational overhead. Using these datasets, we demonstrated that CDR-Agent not only selects relevant CDRs efficiently, but makes cautious yet effective imaging decisions by minimizing unnecessary interventions while successfully identifying most positively diagnosed cases, outperforming traditional LLM prompting approaches. Code for our work can be found at: https://github.com/zhenxianglance/medagent-cdr-agent
@@ 1 Introduction
Clinical Decision Rules (CDRs) are standardized tools designed to assist clinicians by combining signs, symptoms, and clinical features into decision trees, enabling accurate and consistent, evidence-based bedside decisions (Pines & Carpenter, 2023). Currently, over 700 CDRs span nearly all medical specialties, covering diverse clinical scenarios (Obra et al., 2025; Heerink et al., 2023). These rules utilize patient data to generate composite scores or assessments that stratify patient risk for disease onset, progression, or clinical outcomes, helping clinicians deliver expert-level care regardless of their level of experience (Holmes & Kuppermann, 2014; Chan et al., 2020).
@@
However, across healthcare settings, clinicians face increasing time pressure and cognitive burden when evaluating patients. Whether in outpatient clinics, inpatient wards, or telemedicine encounters, clinicians must rapidly assess complex presentations with limited time and incomplete information. In these environments, tools that can surface the right decision aid at the right time can help mitigate errors and democratize care quality. This challenge is even more pronounced in high-stakes trauma care, where rapid and accurate decisions are essential to avoid the harms associated with missed critical injuries or unnecessary imaging and interventions (Neumann et al., 2024). Although trauma care has been regionalized to concentrate expertise, injured patients often initially present to emergency departments (EDs) without trauma specialization, further emphasizing the need for universally applicable decision-making aids like CDRs (Deshormes et al., 2024; Perry & Stiell, 2006).
@@
Despite the importance of CDRs, clinicians often face significant challenges recalling and applying them appropriately due to the sheer number of rules tailored to specific clinical conditions, organ systems, and patient populations (Kharel et al., 2023). This challenge is not confined to emergency care – clinicians across healthcare settings routinely make rapid decisions under time constraints, often with incomplete data. Whether in prehospital, outpatient, or inpatient medicine, clinicians must balance efficiency with accuracy, making real-time access to relevant CDRs critically important.
@@
Recent advancements have highlighted the integration of Large Language Models (LLMs) into Clinical Decision Support Systems (CDSS), significantly enhancing clinical decision-making processes. LLMs augmented with external medical knowledge, such as literature databases and clinical guidelines, demonstrate superior performance compared to standalone models in various clinical tasks, including COVID-19 outpatient care, medication prescriptions, and diagnostic accuracy (Wang et al., 2024; Ong et al., 2024; Lammert et al., 2024). Additionally, explanations generated by LLMs based on patient notes have been shown to improve clinician agreement rates, emphasizing their potential utility in clinical contexts (Umerenkov et al., 2023). Nevertheless, existing approaches often suffer from limited generalizability due to their application within narrow, specialized healthcare settings (Panicker & George, 2023). Particularly relevant to our work is the study by Zakka et al., which introduced an LLM-based agent with external internet access and an extensive database of clinical calculators derived from MDCalc, coupled with a ClinicalQA dataset, achieving substantial performance improvements over plain LLMs (Zakka et al., 2024).
@@
Compared to existing research focused on automated diagnosis (Abbasian et al., 2024; Shi et al., 2024; McDuff et al., 2023; Tu et al., 2024; Levra et al., 2025), knowledge-intensive medical question answering (Singhal et al., 2023; Saab et al., 2024; Wang et al., 2023), and medication management (Ong et al., 2024), emergency department decision-making uniquely prioritizes accuracy and efficiency amidst challenges posed by incomplete patient information. Prior studies that augment LLMs with external medical knowledge typically employ retrieval-augmented generation (RAG) methods, which select relevant resources to contextualize the LLM’s responses. Although these methods improve accuracy and efficiency relative to standalone models, this alone is insufficient in emergency department scenarios. Existing RAG approaches commonly lack transparency regarding how the LLM integrates retrieved information or verifies the appropriateness of selected resources, limiting interpretability. This absence of clear rationale undermines clinical trust, as emergency clinicians rely heavily on transparent, interpretable decision-making to manage rapid, accurate patient care amidst incomplete and evolving clinical information.
@@
Due to the lack of public benchmarks for validating our CDR-Agent, we evaluate its effectiveness and efficiency on two curated ED datasets: CDR-Bench and a synthetic dataset derived from Pediatric Emergency Care Applied Research Network (PECARN) (Holmes et al., 2013; Kuppermann et al., 2009). The synthetic dataset provides a controlled evaluation setting, where each note is generated from PECARN tabular data and contains a single CDR ground truth label with no missing values. This ensures a structured assessment of model performance in unambiguous clinical scenarios. CDR-Bench consists of real-world clinical notes from various public datasets, annotated with clinician-labeled CDRs. These notes exhibit diverse writing styles, contain noise, and often have missing values, making CDR-Bench a challenging real-world robustness test. CDR-Bench features an adaptive number of CDRs per note, better reflecting real-world clinical decision-making complexity. Together, these datasets offer a comprehensive evaluation of CDR-Agent’s ability to interpret clinical narratives and apply trauma-related CDRs accurately. Both datasets are released with code and detailed documentation, ensuring reproducibility and facilitating further research in LLM-driven clinical decision support.
@@
In summary, we propose CDR-Agent, the first automated, LLM-based system that assists with expert-level decision-making using CDRs in ED scenarios. We also created two datasets to facilitate research on ED decision-making involving CDR. Finally, our in-depth case studies underscore the potential to foster collaboration between clinicians and technicians in this field.
@@ 4 Dataset Construction
CDR-Bench is constructed using real-world clinical notes from three well-established public datasets: MIMIC-IV (Johnson et al., 2023), MedQA (Jin et al., 2020), and Augmented Clinical Notes (ACN) (Bonnet, 2023), each featuring diverse writing styles. To ensure a balanced dataset with both CDR-applicable and non-applicable cases, we filter for notes containing trauma-related keywords (e.g., fracture, injury, trauma). This step prevents an overwhelming number of notes without relevant CDRs, given the large dataset sizes.
@@ 7 Related Work
In addition to the related work mentioned in the introduction, there are several other studies also highlight the advancements in clinical decision support systems (CDSS) using LLMs. For example, LLMs enhanced with clinical guidelines show improved performance in tasks like COVID-19 outpatient treatment (Oniani et al., 2022; Wang et al., 2024). For another example, LLM-generated explanations increase doctors’ agreement rates with the diagnoses (Umerenkov et al., 2023). Moreover, various studies employ retrieval augmented generation (RAG) with LLMs or integrated non-knowledge-based models like fine-tuned BERT for medical predictions (Ong et al., 2024; Levra et al., 2025). Differing from these studies, our research specifically focuses on a unique clinical decision-making scenario, trauma-related CDR selection and execution in emergency rooms, where both the accuracy and efficiency of the decision-making process are critical.

### v2

@@ Abstract
Clinical decision-making in emergency departments (EDs) is complex, rapid, and high-stakes. Clinical Decision Rules (CDRs) are standardized evidence-based tools that combine signs, symptoms, and clinical variables into decision trees to make consistent and accurate diagnoses. CDR usage is often hindered by the clinician’s cognitive load, limiting their ability to quickly recall and apply the appropriate rules. Moreover, current AI systems for clinical support struggle to consistently follow medical knowledge, adapt to safety policies, and coordinate complex decision-making. We introduce CDR-Agent, an LLM-based system that autonomously identifies and applies the appropriate CDRs from unstructured clinical notes while maintaining efficiency, robustness, and structured knowledge access. To validate CDR-Agent, we curated two novel ED datasets: synthetic and CDR-Bench, although CDR-Agent is applicable to non ED clinics. CDR-Agent improves CDR selection accuracy by 56.3% on the synthetic dataset and 8.7% on CDR-Bench compared with a standalone LLM baseline, while also reducing computational overhead. Using these datasets, we demonstrated that CDR-Agent not only selects relevant CDRs efficiently, but makes cautious yet effective imaging decisions by minimizing unnecessary interventions while successfully identifying most positively diagnosed cases, outperforming traditional LLM prompting approaches.
@@ 1 Introduction
Clinical Decision Rules (CDRs) are standardized tools designed to assist clinicians by combining signs, symptoms, and clinical features into decision trees, enabling accurate and consistent, evidence-based bedside decisions (Pines2023). Currently, over 700 CDRs span nearly all medical specialties, covering diverse clinical scenarios (Obra2025; Heerink2023). These rules utilize patient data to generate composite scores or assessments that stratify patient risk for disease onset, progression, or clinical outcomes, helping clinicians deliver expert-level care regardless of their level of experience (Holmes2014; Chan2020).
@@
However, across healthcare settings, clinicians face increasing time pressure and cognitive burden when evaluating patients. Whether in outpatient clinics, inpatient wards, or telemedicine encounters, clinicians must rapidly assess complex presentations with limited time and incomplete information. In these environments, tools that can surface the right decision aid at the right time can help mitigate errors and democratize care quality. This challenge is even more pronounced in high-stakes trauma care, where rapid and accurate decisions are essential to avoid the harms associated with missed critical injuries or unnecessary imaging and interventions (Neumann2024). Although trauma care has been regionalized to concentrate expertise, injured patients often initially present to emergency departments (EDs) without trauma specialization, further emphasizing the need for universally applicable decision-making aids like CDRs (Deshormes2024; Perry2006).
@@
Despite the importance of CDRs, clinicians often face significant challenges recalling and applying them appropriately due to the sheer number of rules tailored to specific clinical conditions, organ systems, and patient populations (Kharel2023). This challenge is not confined to emergency care – clinicians across healthcare settings routinely make rapid decisions under time constraints, often with incomplete data. Whether in prehospital, outpatient, or inpatient medicine, clinicians must balance efficiency with accuracy, making real-time access to relevant CDRs critically important.
@@
Recent advancements have highlighted the integration of Large Language Models (LLMs) into Clinical Decision Support Systems (CDSS), significantly enhancing clinical decision-making processes. LLMs augmented with external medical knowledge, such as literature databases and clinical guidelines, demonstrate superior performance compared to standalone models in various clinical tasks, including COVID-19 outpatient care, medication prescriptions, and diagnostic accuracy (Wang2024; Ong2024; Lammert2024). Additionally, explanations generated by LLMs based on patient notes have been shown to improve clinician agreement rates, emphasizing their potential utility in clinical contexts (Umerenkov2023). Nevertheless, existing approaches often suffer from limited generalizability due to their application within narrow, specialized healthcare settings (Panicker2023). Particularly relevant to our work is the study by Zakka et al., which introduced an LLM-based agent with external internet access and an extensive database of clinical calculators derived from MDCalc, coupled with a ClinicalQA dataset, achieving substantial performance improvements over plain LLMs (Zakka2024).
@@
Compared to existing research focused on automated diagnosis (abbasian2024conversational; shi2024ehragent; mcduff2023differential; tu2024diagnosticai; LEVRA2025113), knowledge-intensive medical question answering (SinghalKaran2023Llme; saab2024medgemini; Wang2023), clinical data analysis (Xiao2025.06.01.25328537), and medication management (Ong2024), emergency department decision-making uniquely prioritizes accuracy and efficiency amidst challenges posed by incomplete patient information. Prior studies that augment LLMs with external medical knowledge typically employ retrieval-augmented generation (RAG) methods, which select relevant resources to contextualize the LLM’s responses. Although these methods improve accuracy and efficiency relative to standalone models, this alone is insufficient in emergency department scenarios. Existing RAG approaches commonly lack transparency regarding how the LLM integrates retrieved information or verifies the appropriateness of selected resources, limiting interpretability. This absence of clear rationale undermines clinical trust, as emergency clinicians rely heavily on transparent, interpretable decision-making to manage rapid, accurate patient care amidst incomplete and evolving clinical information.
@@
Due to the lack of public benchmarks for validating our CDR-Agent, we evaluate its effectiveness and efficiency on two curated ED datasets: CDR-Bench and a synthetic dataset derived from Pediatric Emergency Care Applied Research Network (PECARN) (Holmes2013LowRiskAbdominalInjury; Kuppermann2009LowRiskBrainInjury). The synthetic dataset provides a controlled evaluation setting, where each note is generated from PECARN tabular data and contains a single CDR ground truth label with no missing values. This ensures a structured assessment of model performance in unambiguous clinical scenarios. CDR-Bench consists of real-world clinical notes from various public datasets, annotated with clinician-labeled CDRs. These notes exhibit diverse writing styles, contain noise, and often have missing values, making CDR-Bench a challenging real-world robustness test. CDR-Bench features an adaptive number of CDRs per note, better reflecting real-world clinical decision-making complexity. Together, these datasets offer a comprehensive evaluation of CDR-Agent’s ability to interpret clinical narratives and apply trauma-related CDRs accurately. Both datasets are released with code and detailed documentation, ensuring reproducibility and facilitating further research in LLM-driven clinical decision support.
@@
In summary, we propose CDR-Agent, the first automated, LLM-based system that assists with expert-level decision-making using CDRs in ED scenarios. We also created two datasets to facilitate research on ED decision-making involving CDR. Finally, our in-depth case studies underscore the potential to foster collaboration between clinicians and technicians in this field. Code for our work can be found at: https://github.com/zhenxianglance/medagent-cdr-agent
@@ 4 Dataset Construction
CDR-Bench is constructed using real-world clinical notes from three well-established public datasets: MIMIC-IV (johnson2023mimic), MedQA (jin2020disease-medqa), and Augmented Clinical Notes (ACN) (augmented_clinical_notes-acn), each featuring diverse writing styles. To ensure a balanced dataset with both CDR-applicable and non-applicable cases, we filter for notes containing trauma-related keywords (e.g., fracture, injury, trauma). This step prevents an overwhelming number of notes without relevant CDRs, given the large dataset sizes.
@@ 7 Related Work
In addition to the related work mentioned in the introduction, there are several other studies also highlight the advancements in clinical decision support systems (CDSS) using LLMs. For example, LLMs enhanced with clinical guidelines show improved performance in tasks like COVID-19 outpatient treatment (oniani2022enhancing; Wang2024). For another example, LLM-generated explanations increase doctors’ agreement rates with the diagnoses (Umerenkov2023). Moreover, various studies employ retrieval augmented generation (RAG) with LLMs or integrated non-knowledge-based models like fine-tuned BERT for medical predictions (Ong2024; LEVRA2025113). Differing from these studies, our research specifically focuses on a unique clinical decision-making scenario, trauma-related CDR selection and execution in emergency rooms, where both the accuracy and efficiency of the decision-making process are critical.